the perimeter of a rectangular garden is 72 meter the length is 6 times longer than twice the width find dimensions of garden
Answer by JulietG(1812)

P = 2L + 2W
P = 72
L = 12W (that's an odd way to put it -- 6x 2x)
P = 2(12W) + 2W
P = 24W + 2W
72 = 26W
W = 2.769

I *think* the problem should have read "6m longer than twice the width"
In that case:
P = 2L + 2W
P = 72
L = 2W + 6
72 = 2(2W+6) + 2W
72 = 4W+12+2W
72 = 6W+12
60 = 6W
W = 10
If the width is 10, then the length is twice that plus six, or 26
26+26+10+10 = 72